OpenStudy (anonymous):

square root 10 is between what two whole numbers ?

OpenStudy (anonymous):

3 and 4

OpenStudy (smurfy14):

3^2=9 4^2=16

OpenStudy (anonymous):

3^2=9 4^2=16 9<10<16

OpenStudy (anonymous):

okay thank you